# KAAE: Numerical Reasoning for Knowledge Graphs via Knowledge-aware Attributes Learning

### v1

@@
Abstract.
Numerical reasoning is pivotal in various artificial intelligence applications, such as natural language processing and recommender systems, where it involves using entities, relations, and attribute values (e.g., weight, length) to infer new factual relations (e.g., the Nile is longer than the Amazon). However, existing approaches encounter two critical challenges in modeling: (1) semantic relevance—the challenge of insufficiently capturing the necessary contextual interactions among entities, relations, and numerical attributes, often resulting in suboptimal inference; and (2) semantic ambiguity—the difficulty in accurately distinguishing ordinal relationships during numerical reasoning, which compromises the generation of high-quality samples and limits the effectiveness of contrastive learning. To address these challenges, we propose the novel Knowledge-Aware Attributes Embedding model (KAAE) for knowledge graph embeddings in numerical reasoning. Specifically, to overcome the challenge of semantic relevance, we introduce a Mixture-of-Experts-Knowledge-Aware (MoEKA) Encoder, designed to integrate the semantics of entities, relations, and numerical attributes into a joint semantic space. To tackle semantic ambiguity, we implement a new ordinal knowledge contrastive learning (OKCL) strategy that generates high-quality ordinal samples from the original data with the aid of ordinal relations, capturing fine-grained semantic nuances essential for accurate numerical reasoning. Experiments on three public benchmark datasets demonstrate the superior performance of KAAE across various attribute value distributions.
Numerical Reasoning, Knowledge Aware, Knowledge Graph
ccs: Computing methodologies Knowledge representation and reasoning
Figure 1. An example of KG-based numerical reasoning. Dashed lines indicate facts that are not relevant to the query, solid lines and underlined attributes indicate relevant facts.
1. Introduction
Knowledge Graphs (KGs) serve as structured representations of real-world facts and knowledge, playing a pivotal role in powering a wide array of downstream applications, such as recommender systems (RS) (Cao et al., 2022a; Li et al., 2020; Huang et al., 2019), natural language processing (NLP) (Guo et al., 2023; Lu et al., 2023; Sun et al., 2023) and multimodal tasks(Wang et al., 2023; Li et al., 2023; Ni et al., 2023). Among them, numerical reasoning within KGs emerges as a crucial capability, aimed at identifying and inferring missing numerical relations or hierarchical orders among entities.
Existing methods typically integrate additional attributes of entities into traditional knowledge graph embedding (KGE) models, leveraging entities with rich attribute information for numerical reasoning tasks. Others embed numerical entities directly into a vector space or use Graph Neural Networks (GNNs) to enhance KGE performance. However, these approaches often face critical limitations:
@@ 1. Introduction
    Semantic relevance: Most current models fail to capture the intricate interactions among entities, relations, and numerical attributes sufficiently. This deficiency results in suboptimal inference outcomes, as these models do not account for how the significance of numeric attributes changes depending on the relational context. For example, in addressing the query (LeBron James, Taller_Than, ?), the height attribute is crucial, whereas for (LeBron James, Younger_Than, ?), age becomes relevant. This selective relevance is often overlooked, leading to the underutilization of semantic details inherent in the numeric attributes, as depicted in Figure 1.
@@
    Semantic ambiguity: Existing models struggle with accurately distinguishing ordinal relationships between numerical values, which are vital for high-quality inference. Numerical values close in magnitude (e.g., 183cm and 184cm) often carry similar semantic meanings but are prone to erroneous ordinal interpretations under current systems. Such ambiguities can significantly impair the model’s ability to perform effective numerical reasoning.
Given these considerations, we pose the following question: Can we develop a novel KGE model that can conduct knowledge-aware attributes learning and exploit latent numerical information within a unified framework?
To address these challenges, we introduce a new model termed KAAE (Knowledge-Aware Attributes Embedding), crafted for numerical reasoning via knowledge graphs. Specifically, to tackle (C1), we utilize the Mixture-of-Experts (MoE) framework (Bian et al., 2023) to integrate semantic-guided knowledge experts, constructing expert networks for entities across various relational contexts. Additionally, we implement a joint interaction framework that leverages appropriate contexts and captures the intricate connections between entity-relation pairs and attributes within a unified vector space. Employing this knowledge-aware mechanism, our model discerns the significance and relevance of each attribute, effectively synthesizing the latent semantics among them. To address (C2), diverging from prior work that primarily focuses on positive and negative samples, we introduce a novel ordinal knowledge contrastive learning method to generate high-quality ordinal samples. We optimize these ordinal samples based on cosine similarity, selecting the top-k samples with the highest similarity to enhance the learning of ordinal relations. This approach enables the model to capture the continuous features of numerical attributes and discern the subtle differences in boundary values, thus enhancing numerical reasoning. Furthermore, we provide theoretical guarantees for the effectiveness of our model. Extensive experiments show our model can achieve the state-of-the -art performance.
In summary, our contributions are three-fold:
  • To conduct knowledge-aware attribute learning, we introduce a novel approach, KAAE, from semantic interaction and perceptual perspectives. To our knowledge, this is the first model to address semantic relevance by integrating entity relations and numerical attributes within a unified framework for numerical reasoning tasks.
  • To capture the fine-grained semantics of continuous features for numerical attributes, we propose a novel ordinal knowledge contrastive learning method that generates high-quality ordinal samples. This approach enhances the model’s ability to perform nuanced numerical reasoning by exploiting ordinal relationships within the data.
  • Extensive experiments show that KAAE achieves state-of-the-art performance on three real-world datasets, and demonstrate its robustness, reasonability, and interpretability.
2. Related Work
@@
Knowledge Graph Embedding(KGE). KGE amis to capture the latent representations of entities and relations in KGs. TransE(Bordes et al., 2013) and its extended models(Wang et al., 2014; Zhang et al., 2021; Ji et al., 2015), focus on treating relation as a ”translation” from the head entity to the tail entity. DistMult(Dong et al., 2014) represent each relation with a diagonal matrix, ComplEx(Trouillon et al., 2016) extends DistMult to complex space. ConvE(Dettmers et al., 2018) extracts deep features of head entity and relation based on 2D convolution. ConE(Bai et al., 2021), MuRP(Balazevic et al., 2019a), GIE(Cao et al., 2022b) and AttH(Chami et al., 2020) embed KGs into hyperbolic spaces to model hierarchical relations. Further, models(Schlichtkrull et al., 2018; Shang et al., 2019; Vashishth et al., 2020) of GNNs have been proposed to model higher-order connectivity in knowledge graphs. However, these models do not consider numerical values, making it difficult to accomplish numerical reasoning tasks.
@@
Knowledge Graph Embedding with Numerical Attributes. To improve the performance of KGE, several KGE models have been proposed to incorporate auxiliary information about the entity such as literals and numerical attributes (e.g., age, weight, and born_year). For example, KBLRN(García-Durán and Niepert, 2018) accomplish the KGE task based on numerical values, considering numerical differences between different entities. MT-KGNN(Tay et al., 2017) and TransEA(Wu and Wang, 2018) are models for multi-task learning to predict both numeric attribute values and entity/relation embeddings. LiteralE(Kristiadi et al., 2019) combines the entity embedding with numeric attributes based on a learnable gating function. Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023) are also used to predict numerical attributes in KGs. Moreover, RAKGE(Kim et al., 2023) learns the representation for attribution with relations. However, these methods do not consider the joint information among entities, attributes, and relations. On the contrary, our KAAE model can exploit joint complex interactions among entities, relations and numerical attributes to fully mine the semantic information useful for numerical reasoning tasks.
@@
Contrastive Learning. Contrastive learning (CL) is an effective self-supervised learning method by pulling semantically close neighbors together while pushing non-neighbors away, and some models have attempted to utilize contrastive learning to improve representations of entities/graphs. SimGCL(Yu et al., 2022) adds uniform noise to the entity representation, which is an expansion-free CL method. HeCO(Wang et al., 2021b) employs contrast learning in network schema and meta-path, capturing both local and global features of entities. SLiCE(Wang et al., 2021a) makes use of mutual attraction between closed nodes to learn subgraph representations. RAKGE(Kim et al., 2023) interacts with training samples with head entities to generate samples. However, none of the above methods can guarantee effectiveness in numerical reasoning tasks. To improve contrast learning, we optimize the sampling process with ordinal relations in KGs and generate high-quality ordinal positive/negative samples based on the technique of similarity sampling, thus achieving significant performance improvements.
3. Methodology
@@
In this section, we present a new KG-based numerical reasoning framework called KAAE from the semantic joint interaction perspective. Specifically, there are two main components: Mixture-of-Experts-Knowledge-Aware Encoder (MoEKA-Encoder) and ordinal knowledge contrastive learning (OKCL) strategy. The overall framework of KAAE is illustrated in Figure 2.
@@
Input Knowledge Graph             MoEKA Encoder          Ordinal Knowledge Contrastive Loss
Figure 2. Overview of our KAAE model. The MoEKA Encoder encodes each entity with the relation and attributes to a unified representation thus tackling the semantic relevance. Afterward, to capture the fine-grained semantic information, we utilize a new knowledge contrastive learning method to generate high-quality ordinal samples thus tackling semantic ambiguity. In this way, we can learn the nuanced differences in attributes and perform better numerical reasoning tasks.
3.1. Knowledge-Aware Learning
To tackle the semantic relevance of KGs, we propose a MoE-knowledge-aware encoder consisting of four parts: the numeric value embedding layer, a mixture of knowledge experts, the knowledge perceptual attention layer, and the knowledge aggregation layer.
@@ 3.1. Knowledge-Aware Learning
Mixture of Knowledge Experts. To better learn entity embeddings from different perspectives with relation contexts and numerical attributes, we first employ a module called Mixture of Knowledge Experts to build expert networks, where each perspective corresponds to an expert network. First, the entity $i\in\mathcal{E}$ has a raw feature $\tilde{e}_{i}$. We then learn the multi-perspective embeddings $\mathcal{H}^{\tilde{e}_{i}}_{1},\mathcal{H}^{\tilde{e}_{i}}_{2},\ldots,\mathcal{H}^{\tilde{e}_{i}}_{K}$ for the entity $i$ by establishing $K$ knowledge experts denoted as $\mathcal{W}_{1},\mathcal{W}_{2},\ldots,\mathcal{W}_{K}$. This process can be represented as $\mathcal{H}^{\tilde{e}_{i}}_{k}=\mathcal{W}_{k}(\tilde{e}_{i})$. Next, we design a semantic-guided gated fusion network (SGN) to facilitate the fusion of inter-perspective entity embeddings with relation guidance.
(1) $\bm{e}_{i}=\sum\nolimits_{k=1}^{K}\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)\mathcal{H}^{\tilde{e}_{i}}_{k}$
where $\bm{e_{i}}\in\mathbb{R}^{d_{emb}}$ is the output entity embedding of entity i for relation $r$, $\mathcal{F}_{k}$ is the weight for each expert calculated by the GFN:
(2) $\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)=\frac{exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{k})+\psi_{k})/\rho(\epsilon_{r}))}{\sum\nolimits_{j=1}^{K}exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{j})+\psi_{j})/\rho(\epsilon_{r}))},$
where $\psi_{k}\sim\mathcal{N}(0,\mathcal{U^{\prime}}(\mathcal{H}^{\tilde{e}_{i}}_{k}))$, $\mathcal{U}$ and $\mathcal{U^{\prime}}$ are two projection layers, and $\psi_{k}$ is tunable Gaussian noise used to balance the weights for each expert and enhance the model’s robustness. Additionally, we add a relation-aware temperature $\epsilon_{r}$ with a sigmoid function $\rho$ to limit the temperature within the range $(0,1)$. Our aim is to obtain an entity embedding within the relational context of the current prediction before making the final decision. This approach introduces the relational context to guide embedding learning within the experts, enabling them to extract relation-aware embeddings. Additionally, the experts will be differentiated to adapt to different relational contexts through the design of SGN. This allows us to learn dynamic embeddings of entities that change according to different relational contexts.
Knowledge Perceptual Attention. In the previous sections, we obtained the entity representations within the relational context. Based on this, we aim to learn the knowledge-aware representations via the joint semantic interactions among entities, relations, and attributes. In this way, we propose a knowledge perceptual attention mechanism for entity-relation attributes using multi-head self-attention(Vaswani et al., 2017) to capture the importance of numerical attributes based on each entity and its associated relations. Specifically, we leverage a single layer perception $f_{e-att}$ and $f_{r-att}$ to project the entity and relation embeddings onto the attribute embedding space, respectively. Thereafter, we mixing the information of entity e and relation r, so as to get the joint projected embedding $\bm{p}_{joint}^{att}\in\mathbb{R}^{d_{att}}$:
(3) $\bm{p}_{joint}^{att}=\sum_{*\in\{e,r\}}\delta_{*}f_{*-att}(\bm{p}_{*})$
@@
Then, we apply the knowledge perceptual attention mechanism to establish the joint interactions between entities, relations, and attributes. Specifically, We apply this mechanism for all attribute fields $m$ = $\cdots,|\mathcal{M}|$ and we repeat this step for all attention heads l = 1,2$\cdots$,$L$, where $L$ is the number of multi-head attentions. The result of the joint interaction of head l is shown as follows:
(4) $\bm{o}_{joint,i}^{(l)}=\sum\nolimits_{m=1}^{|\mathcal{M}|}a_{joint,i,m}^{(l)}(\bm{W}_{agg}^{(l)}\bm{o}_{i}^{m}),$
where the linear transformation matrix $\bm{W}_{agg}^{(l)}\in\mathbb{R}^{d_{sub}\times d_{att}}$ aims to project attribute embedding $\bm{o}_{i}^{m}$ into low-dimensional subspaces to capture the importance of each numeric attribute from the given entity with relation accurately. Under a specific attention head l, we map $\bm{e}_{joint}^{att}$ and $\bm{o}_{i}^{1},\cdots\bm{o}_{i}^{|\mathcal{M}|}\in\mathbb{R}^{d_{att}}$ onto smaller subspaces and capture the attention score. The normalized attention weight $a_{joint,i,m}^{(l)}$ of the attribute embedding $\bm{o}_{i}^{m}$ is formulated as follows:
(5) $a_{joint,i,m}^{(l)}=\frac{exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m}))}{\sum\nolimits_{n=1}^{|\mathcal{M}|}exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{n}))},$
and
(6) $s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})=\sigma\left(\|_{j=1}^{H}\bm{W}_{j}^{(l)}\phi_{j}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})\right)$
@@
[…]
(7) $\bm{e}_{joint,i}^{att}=\sigma(\bm{W}_{ge}\bm{e}_{i}+\bm{W}_{ga}\bm{o}_{joint,i})+\bm{b}$
where $\bm{W}{ge}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ and $\bm{W}_{ga}\in\mathbb{R}^{d_{emb}\times d_{att}}$ are different linear transformations to assign different weights to each entity representation and fused numerical representation by adaptively learning to obtain high quality final embeddings and to focus on the more important information. $\bm{b}$ is a bias. $\sigma$ is a sigmoid function, and $\tau$ is a hyperbolic tangent function. We name $\bm{e}_{joint,i}^{att}\in\mathbb{R}^{d_{emb}}$ as the attribute-enriched vector of entity i.
3.2. Ordinal Knowledge Contrastive Learning
Learning ordinal relations is essential for numerical reasoning tasks. Specifically, for numerical reasoning, data consists of continuous scalar values, resulting in numerical values that are extremely sensitive to minimal differences near boundary values, thus affecting the accuracy of ordinal embedding. Unlike the previous work mainly focusing on generating valid, positive, and negative samples, we turn to a more difficult and significant task that generates high-quality ordinal samples. For example, consider the head entity with a weight attribution of 80kg and the relation is_heavier_than. The ideal learning range consists of generating all possible positive tail entities weighing less than 80kg. It is particularly crucial to distinguish boundary values(e.g., tails between 79 and 81kg) more efficiently.
Ordinal Relation Learning. Given $N$ objects and the ordinal relationship triples $\mathcal{S}=\{(a,b,c)\mid a,b,c\in\mathcal{E},a\neq b\neq c\}$, where $[N]=\{1,\ldots,N\}$ and object $a$ is more similar to object $b$ than it is to object $c$. Given some distance function $d(\cdot,\cdot)$, we aim to learn the representations of objects, denoted as $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$, such that the following objectives hold as much as possible:
(8) $d(\mathbf{x}_{a},\mathbf{x}_{b})<d(\mathbf{x}_{a},\mathbf{x}_{c}),\quad\forall(a,b,c)\in\mathcal{S}.$
@@ 3.2. Ordinal Knowledge Contrastive Learning
(9) +=$\centering\mathcal{E^{+}}=\{\bm{e}_{joint,i}^{att}|i\in\mathcal{P}[h,r]\},\ \centering\mathcal{E^{-}}=\{\bm{e}_{joint,j}^{att}|j\in\mathcal{N}[h,r]\},\@add@centering\@add@centering$
where $\mathcal{P}[h,r]$ represents the set of positive tails containing entities whose weight is less than 80kg (e.g., 65, 70, and 75kg). $\mathcal{N}[h,r]$ represents the set of positive tails containing entities whose weight is more than 80kg (e.g., 85, 90, and 95kg).

### v3

# NumCoKE: Ordinal-Aware Numerical Reasoning over Knowledge Graphs with Mixture-of-Experts and Contrastive Learning

@@
Abstract
Knowledge graphs (KGs) serve as a vital backbone for a wide range of AI applications, including natural language understanding and recommendation. A promising yet underexplored direction is numerical reasoning over KGs, which involves inferring new facts by leveraging not only symbolic triples but also numerical attribute values (e.g., length, weight). However, existing methods fall short in two key aspects: (1) Incomplete semantic integration: Most models struggle to jointly encode entities, relations, and numerical attributes in a unified representation space, limiting their ability to extract relation-aware semantics from numeric information. (2) Ordinal indistinguishability: Due to subtle differences between close values and sampling imbalance, models often fail to capture fine-grained ordinal relationships (e.g., longer, heavier), especially in the presence of hard negatives. To address these challenges, we propose NumCoKE—a numerical reasoning framework for KGs based on Mixture-of-Experts and Ordinal Contrastive Embedding. To overcome (C1), we introduce a Mixture-of-Experts Knowledge-Aware (MoEKA) encoder that jointly aligns symbolic and numeric components into a shared semantic space, while dynamically routing attribute features to relation-specific experts. To handle (C2), we propose Ordinal Knowledge Contrastive Learning (OKCL), which constructs ordinal-aware positive and negative samples using prior knowledge, enabling the model to better discriminate subtle semantic shifts. Extensive experiments on three public KG benchmarks demonstrate that NumCoKE consistently outperforms competitive baselines across diverse attribute distributions, validating its superiority in both semantic integration and ordinal reasoning.
Introduction
Knowledge graphs (KGs) represent structured factual knowledge as triples of entities and relations, and have become a key foundation for various AI applications, including recommendation systems (DBLP:conf/mm/CaoSW0WY22; DBLP:conf/mm/LiXJCH20), natural language processing (DBLP:conf/mm/Guo0ZM0WX23; DBLP:conf/mm/LuDYY23; DBLP:conf/mm/SunYLLLS23), and multimodal tasks (DBLP:conf/mm/WangMCML023; DBLP:conf/mm/LiQZ0TXT23). A crucial yet underexplored capability of KGs is numerical reasoning, the ability to infer facts involving quantitative comparisons or numerical ordering (e.g., “the Nile is longer than the Amazon”). This ability is particularly valuable for information services that require precise numerical judgments, such as fine-grained product recommendations or numerical question answering.
Figure 1: Example of KG-based numerical reasoning. Same colored text indicates strong related relation and attribute.
To enhance numerical reasoning, prior studies have explored integrating attribute values into knowledge graph embedding (KGE) frameworks. Some directly embed numerical entities into continuous spaces (DBLP:conf/kdd/BaiLLYYS23), while others employ graph neural networks (DBLP:conf/iclr/VashishthSNT20) or augment traditional KGE with attribute-aware components (DBLP:conf/kdd/KimKKPJP23). Despite these advances, existing methods face two persistent and intertwined challenges:
@@ Introduction
    Incomplete semantic integration. Current models often fail to effectively capture the joint semantics of entities, relations, and numerical attributes. In practice, the relevance of numerical attributes is highly dependent on the relational context. For instance, given the query (San Francisco, Less_Elders, x), the attribute median_age becomes critical; whereas in (San Francisco, Easier_Commuting, x), the key attribute is commute_time. Models lacking context-aware alignment treat numerical features uniformly, leading to semantic incompleteness and degraded reasoning performance. (see Figure 1).
@@
    Ordinal indistinguishability. Numerical reasoning often requires fine-grained distinction between close values. However, existing models typically learn coarse semantic representations of numerical attributes, making it difficult to capture such subtle ordinal relations. This leads to confusion in tasks that rely on comparative inference, and the ambiguity becomes more pronounced when handling hard negative samples with near-equal values.
To address these challenges, we propose NumCoKE, a novel framework for numerical reasoning on knowledge graphs via Mixture-of-Experts and Ordinal Contrastive Embedding. To tackle (C1), we design a Mixture-of-Experts Knowledge-Aware (MoEKA) encoder that dynamically captures the contextual importance of numeric attributes under different relational settings. Specifically, we construct relation-aware expert modules that encode entities and relations into a shared vector space. Through adaptive routing, numerical attributes are directed to the most relevant experts based on their semantic context. This mechanism enables NumCoKE to selectively amplify semantically important attributes for each relation-entity pair, thereby producing fine-grained and context-sensitive representations. To tackle (C2), we propose an Ordinal Knowledge Contrastive Learning (OKCL) strategy. Instead of relying solely on binary positive-negative sampling, OKCL synthesizes ordinal sample triplets by computing cosine-based similarity among value distributions. We then select top-$k$ ordinal neighbors to serve as fine-grained supervision signals. This strategy preserves the continuity of numerical features and explicitly encourages the model to recognize nuanced ordinal relations between close values (e.g., distinguishing 183cm from 184cm). We further provide a theoretical analysis to justify the consistency and effectiveness of OKCL.
We validate the effectiveness of NumCoKE through extensive experiments on three real-world benchmark datasets covering diverse types of numerical attribute distributions (discrete, continuous, and skewed). The results show that our method consistently outperforms strong baselines across all settings. In summary, our contributions are three-folds:
  • We propose NumCoKE, a novel and efficient model for numerical reasoning over KGs, which for the first time incorporates relation-aware mixture-of-experts encoding to dynamically integrate entity, relation, and attribute semantics.
  • We introduce ordinal knowledge contrastive learning, a new contrastive paradigm that generates ordinally structured samples to better capture fine-grained numerical distinctions, improving the model’s sensitivity to subtle ordinal relationships.
  • We conduct comprehensive experiments on three public datasets. NumCoKE achieves state-of-the-art performance, demonstrating superior accuracy of our model.
Related Work
@@
Knowledge Graph Embedding (KGE). KGE amis to capture the latent representations of entities and relations in KGs. TransE (DBLP:conf/nips/BordesUGWY13) and its extended models (DBLP:journals/jifs/ZhangJWZKQL21; DBLP:conf/coling/LiSZG24), focus on treating relation as a translation from the head entity to the tail entity. DaBR (DBLP:conf/coling/WangLBG25) utilize distance-adaptive translations to learn geometric distance between entities, KGDM (DBLP:conf/aaai/LongZLWLW24) estimates the probabilistic distribution of target entities in prediction through diffusion models. RuIE (DBLP:conf/coling/LiaoDH025) leverage logical rules to enhance reasoning. ConvE (DBLP:conf/aaai/DettmersMS018) extracts deep features of head entity and relation based on 2D convolution. ConE (DBLP:conf/nips/BaiYRL21), MuRP (DBLP:conf/nips/BalazevicAH19), GIE (DBLP:conf/aaai/CaoX0CH22), AttH (DBLP:conf/acl/ChamiWJSRR20) and LorentzKG (DBLP:conf/acl/Fan0CCDY24) embed KGs into hyperbolic spaces to model hierarchical relations. Further, models (DBLP:conf/esws/SchlichtkrullKB18; DBLP:conf/aaai/ShangTHBHZ19; DBLP:conf/iclr/VashishthSNT20) of GNNs have been proposed to model higher-order connectivity in knowledge graphs. However, these models do not consider numerical values, making it difficult to accomplish numerical reasoning tasks.
@@
KGE with Numerical Attributes. To improve the performance of KGE, several KGE models have been proposed to incorporate auxiliary information about the entity such as literals and numerical attributes. For example, KBLRN (DBLP:conf/uai/Garcia-DuranN18) accomplish the KGE task based on numerical values, considering numerical differences between different entities. MT-KGNN (DBLP:conf/cikm/TayTPH17) is a multitask model to predict both numeric attribute values and entity/relation embeddings. LiteralE (DBLP:conf/semweb/KristiadiKL0F19) combines the entity embedding with numeric attributes based on a learnable gating function. Deterministic stand-alone value representation methods including NEKG (DBLP:conf/emnlp/DuanYT21) and NRN (DBLP:conf/kdd/BaiLLYYS23) are also used to predict numerical attributes in KGs. However, these models do not exploit joint interactions among entities, relations and numerical attributes to mine the useful semantic information.
@@
Contrastive Learning (CL). CL is a self-supervised learning method by pulling semantically close neighbors together while pushing non-neighbors away, which can improving the representations of entities/graphs. SimGCL (DBLP:conf/sigir/YuY00CN22) adds uniform noise to the entity representation, which is an expansion-free CL method. HeCO (DBLP:conf/kdd/WangLHS21) employs contrast learning in network schema and meta-path, capturing both local and global features of entities. SLiCE (DBLP:conf/www/WangAHCR21) makes use of mutual attraction between closed nodes to learn subgraph representations. RAKGE (DBLP:conf/kdd/KimKKPJP23) interacts with training samples with head entities to generate samples. RRNE (DBLP:conf/pakdd/JeongJKKKP25) selects hard samples from subgraph. However, none of the above methods can guarantee effective samples for numerical reasoning tasks.
Methodology
@@
In this section, we present a new numerical reasoning framework for KG termed NumCoKE. Specifically, there are two main components: Mixture-of-Experts-Knowledge-Aware Encoder (MoEKA-Encoder) and ordinal knowledge contrastive learning (OKCL) strategy. The overall framework of NumCoKE is illustrated in Figure 2.
@@
Figure 2: The Overview of our model. (a) The MoEKA Encoder encodes each entity with the relation and attributes to a unified, elaborate semantic representation. (b) To capture the fine-grained semantic information, we utilize a new knowledge contrastive learning method to generate high-quality ordinal samples to learn the nuances in attributes and distinguish similar semantics.
Knowledge-Aware Learning
@@
Mixture of Knowledge Experts (MOE). To better learn entity embeddings from different perspectives with relation contexts and numerical attributes, we first employ a module called Mixture of Knowledge Experts to build expert networks, where each perspective corresponds to an expert network. First, the entity $i\in\mathcal{E}$ has a raw feature $\tilde{e}_{i}$. We then learn the multi-perspective embeddings $\mathcal{H}^{\tilde{e}_{i}}_{1},\mathcal{H}^{\tilde{e}_{i}}_{2},\ldots,\mathcal{H}^{\tilde{e}_{i}}_{K}$ for the entity $i$ by establishing $K$ knowledge experts denoted as $\mathcal{W}_{1},\mathcal{W}_{2},\ldots,\mathcal{W}_{K}$. This process can be represented as $\mathcal{H}^{\tilde{e}_{i}}_{k}=\mathcal{W}_{k}(\tilde{e}_{i})$. Next, we design a semantic-guided gated fusion network (SGN) to facilitate the fusion of inter-perspective entity embeddings with relation guidance.
$\bm{e}_{i}=\sum\nolimits_{k=1}^{K}\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)\mathcal{H}^{\tilde{e}_{i}}_{k}$ (1)
where $\bm{e_{i}}\in\mathbb{R}^{d_{emb}}$ is the output entity embedding of entity i for relation $r$, $\mathcal{F}_{k}$ is the weight for each expert:
$\mathcal{F}_{k}(\mathcal{H}^{\tilde{e}_{i}}_{k},r)=\frac{exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{k})+\psi_{k})/\rho(\epsilon_{r}))}{\sum\nolimits_{j=1}^{K}exp((\mathcal{U}(\mathcal{H}^{\tilde{e}_{i}}_{j})+\psi_{j})/\rho(\epsilon_{r}))}$ (2)
where $\psi_{k}\sim\mathcal{N}(0,\mathcal{U^{\prime}}(\mathcal{H}^{\tilde{e}_{i}}_{k}))$, $\mathcal{U}$ and $\mathcal{U^{\prime}}$ are two projection layers, and $\psi_{k}$ is tunable Gaussian noise (DBLP:conf/cikm/BianPZWWW23) used to balance the weights for each expert and enhance the model’s robustness. Additionally, we add a relation-aware temperature $\epsilon_{r}$ with a sigmoid function $\rho$ to limit the temperature within the range $(0,1)$. Our aim is to obtain an entity embedding within the relational context of the current prediction before making the final decision.
Knowledge Perceptual Attention. Based on the representations above, we then learn the knowledge-aware representations via the joint semantic interactions among entities, relations, and attributes. In this way, we propose a knowledge perceptual attention mechanism for entity-relation attributes. Specifically, we leverage a single-layer perception $f_{e-att}$ and $f_{r-att}$ to project the entity and relation embeddings onto the attribute embedding subspace, respectively. Thereafter, we mixing the information of entity e and relation r, so as to get the joint projected embedding $\bm{p}_{joint}^{att}\in\mathbb{R}^{d_{att}}$:
$\bm{p}_{joint}^{att}=\sum_{*\in\{e,r\}}\delta_{*}f_{*-att}(\bm{p}_{*})$ (3)
@@
$\bm{o}_{joint,i}^{(l)}=\sum\nolimits_{m=1}^{|\mathcal{M}|}a_{joint,i,m}^{(l)}(\bm{W}_{agg}^{(l)}\bm{o}_{i}^{m})$ (4)
where the linear transformation matrix $\bm{W}_{agg}^{(l)}\in\mathbb{R}^{d_{sub}\times d_{att}}$ aims to project attribute embedding $\bm{o}_{i}^{m}$ into low-dimensional subspaces to capture the importance of each numeric attribute from the given entity with relation accurately. Under a specific attention head l, we map $\bm{e}_{joint}^{att}$ and $\bm{o}_{i}^{1},\cdots\bm{o}_{i}^{|\mathcal{M}|}\in\mathbb{R}^{d_{att}}$ onto smaller spaces and capture the attention score. The normalized attention weight $a_{joint,i,m}^{(l)}$ of the attribute embedding $\bm{o}_{i}^{m}$ is formulated as follows:
$a_{joint,i,m}^{(l)}=\frac{exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m}))}{\sum\nolimits_{n=1}^{|\mathcal{M}|}exp(s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{n}))}$ (5)
$s^{(l)}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})=\sigma\left(\|_{j=1}^{H}\bm{W}_{j}^{(l)}\phi_{j}(\bm{p}_{joint}^{att},\bm{o}_{i}^{m})\right)$ (6)
where $\sigma$ denotes the LeakyReLU activation function, $\bm{W}_{j}^{(l)}$ denotes the weighted matrix corresponding to the $i$-th operator in the $l$-th layer, $\phi_{j}((\bm{p}_{joint}^{att},\bm{o}_{i}^{m}))$ denotes the fusion operation, which can be model by MLP. $\|$ here denotes the multi-head mechanism.
Knowledge-aware Aggregation. The final attribute embedding for entity i related to the joint perceptual vector is as follows: $\bm{o}_{joint,i}=\bm{o}_{joint,i}^{(1)}\|\bm{o}_{joint,i}^{(2)}\|\cdots\|\bm{o}_{joint,i}^{(L)}$, where $\bm{o}_{joint,i}\in\mathbb{R}^{d_{att}}$ as the attribute vector of entity $i$, which is aware of the entity and the relation. Symbol $\|$ represents a concatenation operator. To balance the joint knowledge-aware attribute vectors and the embedded entities, we have:
$\bm{e}_{joint,i}^{att}=\sigma(\bm{W}_{ge}\bm{e}_{i}+\bm{W}_{ga}\bm{o}_{joint,i})+\bm{b}$ (7)
where $\bm{W}{ge}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ and $\bm{W}_{ga}\in\mathbb{R}^{d_{emb}\times d_{att}}$ are different linear transformations to assign different weights to each entity representation and fused numerical representation by adaptively learning to obtain high quality final embeddings and to focus on the more important information. $\bm{b}$ is a bias. $\sigma$ is a sigmoid function. We name $\bm{e}_{joint,i}^{att}\in\mathbb{R}^{d_{emb}}$ as the attribute-enriched vector of entity i.
Ordinal Knowledge Contrastive Learning
Learning ordinal relations is essential for numerical reasoning tasks. Unlike the previous work (DBLP:conf/kdd/KimKKPJP23) mainly focusing on generating valid, positive, and negative samples, we turn to a more difficult and significant task that generates high-quality ordinal samples.
Ordinal Relation Learning. Given $N$ objects and the ordinal relationship triples $\mathcal{S}=\{(a,b,c)\mid a,b,c\in\mathcal{E},a\neq b\neq c\}$, where $N]=\{1,\ldots,N\}$ and object $a$ is more similar to object $b$ than it is to object $c$. Given some distance function $d(\cdot,\cdot)$, we aim to learn the representations of objects, denoted as $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}\}$, such that the following objectives hold as much as possible:
$d(\mathbf{x}_{a},\mathbf{x}_{b})<d(\mathbf{x}_{a},\mathbf{x}_{c}),\quad\forall(a,b,c)\in\mathcal{S}$ (8)
@@
$\centering\mathcal{E^{+}}=\{\bm{e}_{joint,i}^{att}|i\in\mathcal{P}[h,r]\},\ \centering\mathcal{E^{-}}=\{\bm{e}_{joint,j}^{att}|j\in\mathcal{N}[h,r]\}\@add@centering\@add@centering$ (9)
where $\mathcal{P}[h,r]$ represents the set of positive tail entities, $\mathcal{N}[h,r]$ represents the set of positive tail entities.